OpenStudy (anonymous):

. . If 1/2x+7=16, what is the value of 2x + 1? A. -5 B. -27 C. -53 D. -137

myininaya (myininaya):

\[\frac{1}{2x+7}=16?\]

myininaya (myininaya):

or is it \[\frac{1}{2x}+7=16\]

myininaya (myininaya):

or is it \[\frac{1}{2}x+7=16\]

myininaya (myininaya):

1st, 2nd, or 3rd?

myininaya (myininaya):

\[\frac{1}{2}x=16-7\] \[\frac{1}{2}x=9\] \[x=2*9=18\] so 2*18+1=37

myininaya (myininaya):

so i you don't mean the third one do you? i'm not playing guessing game anymore

OpenStudy (anonymous):

yes its c... but i dont know how you did it

myininaya (myininaya):

what is the equation you have? you never answer that i assumed it was the third one because base on what you wrote it is that but i didn't get any of those answers so can please tell me of the equations i wrote which one you meant

OpenStudy (anonymous):

how did you get c i know thats the answer

OpenStudy (anonymous):

idk??? how to get the answer i just know its c

myininaya (myininaya):

i know but which of the equations is it suppose be do you see the one's i wrote above?

OpenStudy (anonymous):

arboring, we're having trouble with your typing of the problem. As myininaya points out, we could interpret the expression you typed in several ways, each of which would give a different answer. however, as you say the answer is C. -53, we can work backwards to establish that x = -27. So now we have to work out what expression in x on the LHS of your expression = 16.

OpenStudy (anonymous):

this is a question from a practice thea test i copied it word for word on here, so there is no mistakes in the question.

myininaya (myininaya):

give me the link the the practice test because it seems there are grouping symbols missing